OpenStudy (anonymous):

Fill in the missing value for the blank. 12x^4+24x^3-6x^2(2x^2+blank-1)

OpenStudy (anonymous):

We can replace 'blank' with a variable, y. This will help us think about the problem, as all we really need to do is solve for y. 12x^4 + 24x^3 - 6x^2(2x^2 + y - 1) = 0 12x^4 +24x^3 - 12x^4 -6x^2y - 6x^2 = 0 24x^3 - 6x^2 = 6x^2y Now divide both sides by 6x^2 to isolate y 4x - 1 = y so 'blank' is 4x - 1